Concurrency: Practice and Experience Papers Accepted for Publication — March 8, 1999

(C323) “Dynamic Scheduling of Process Groups”

Authors: K. Y. Wang, D. C. Marinescu, and O. F. Carbunar Current Status: Paper to Publisher (3/4/97). 34 pgs. Short Code: [Wang:97a]

(C340) “Visualization in Network Computing Environments”

Authors: Brad Topol and John T. Stasko Current Status: Paper to Publisher (8/27/97). 42 pgs. Short Code: [Topol:97a]

(C351) “SCIPVM: Parallel Distributed Computing on SCI Workstation Clusters”

Authors: Ivan Zoraja, Hermann Hillwagner, and Vaidy Sunderam Current Status: Paper to Publisher (10/3/97). 22 pgs. Short Code: [Zoraja:97a]

(C353) “Tiling with Limited Resources”

Authors: Pierre-Yves Calland, Jack Dongarra, and Yves Robert Current Status: Paper to Publisher (12/29/97). 18 pgs. Short Code: [Calland:97a]

(C358) “An Adaptive Load Distribution Algorithm for Resolving Burst Workload”

Authors: Qin Lu, and Sau-Ming Lau Current Status: Paper to Publisher (6/23/98). 34 pgs. Short Code: [Lu:98b]

(C365) “A Computing Framework for Integrating Interactive Visualization in HPCC Applications” Authors: G. Cheng, G. C. Fox, T.-H. Lin, and T. Haupt Current Status: Paper to Publisher (6/30/97). 21 pgs. Short Code: [Cheng:96c] (C366) “The European Porting Project—Demonstrating Industrial Relevance of Parallel HPC”

Authors: S. Brindle, A. Colbrook, J. Elliott, D. Green, and F. Wray Current Status: Paper to Publisher (6/30/97). 19 pgs. Short Code: [Brindle:96a]

(C367) “Parallel Computation of Three-Dimensional Nonlinear Magnetostatic Problems”

Authors: David Levine, William Gropp, Kimmo Forsman, and Lauri Kettunen Current Status: Paper to Publisher (6/30/97). 13 pgs. Short Code: [Levine:96a]

(C368) “Composing High Performance Schedulers: A Case Study from Real-Time Simulation” Authors: K. Ghosh, R. M. Fujimoto, and K. Schwan Current Status: Paper to Publisher (3/4/98). 34 pgs. Short Code: [Ghosh:98a]

(C371) “Scheduling Tasks of Multi-Join Queries in a Multiprocessor”

Authors: A. Averbuch, Y. Roditty, and B. Shoham Current Status: Paper to Publisher (6/23/98). 41 pgs. Short Code: [Averbuch:98a]

(C372) “Multiresolution Load Balancing in Curved Space: The Wavelet Representation”

Authors: Aiichiro Nakano Current Status: Paper to Publisher (11/2/98). 20 pgs. Short Code: [Nakano:98a]

(C374) “A Software Design Technique for Client-server Applications”

Authors: Jahangir Karimi Current Status: Paper to Publisher (6/24/98). 27 pgs. Short Code: [Karimi:98a]

(C375) “Evaluating the Communications Capabilities of the Generalized Hypercube Interconnection Network” Authors: Sanjay Krishnamurthy and Sotirios G. Ziavras Current Status: Paper to Publisher (11/2/98). 35 pgs. Short Code: [Ziavras:98a] (C376) “Parallelism and Recursion in Message Passing Libraries: An Efficient Methodology”

Authors: C. Rodriguez, F. de Sande, C. Leon, and L. Garcia Current Status: Paper to Publisher (11/2/980. 18 pgs. Short Code: [Rodriguez:98a]

(C380) “Parallel Execution to Embedded and Iterated Runge-Kutta Methods”

Authors: Thomas Rauber and Gudula Rünger Current Status: Paper to Publisher (11/2/98). 35 pgs. Short Code: [Rauber:98a]

(C381) “Prophet: Automated Scheduling of SPMD Programs in Workstation Networks” Authors: Jon B. Weissman Current Status: Paper to Publisher (11/2/98). 33 pgs. Short Code: [Weissman:98a]

(C382) “COLT_{HPF}, a Coordination Layer for HPF Tasks”

Authors: Salvatore Orlando and Raffaele Perego Current Status: Paper to Publisher (2/16/99). 37 pgs. Short Code: [Orlando:98a]

(C383) “Predicting the Execution Time of Message Passing Models”

Authors: J. L. Roda, C. Rodriguez, D. G. Morales, and F. Almeida Current Status: Paper to Publisher (2/16/99). 31 pgs. Short Code: [Roda:99a]

(C386) “Concurrent Simulation of Neutral Flow in the GEC Reference Cell”

Authors: M. Rieffel, M. Ivanov, S. Shankar, and S. Taylor Current Status: Paper to Publisher (2/16/99). 28 pgs. Short Code: [Rieffel:98a]

(C388) “Usefulness of Adaptive Load Sharing for Parallel Processing on Networks of Workstations” Authors: Sheldon Clark and Sivarama P. Dandamudi Current Status: Paper to Publisher (2/16/99). 33 pgs. Short Code: [Clark:98a] (C389) “A Communication Micro-Kernel for Implementing Object-Oriented Programming on Closely Coupled

Distributed Architectures” Authors: Giullano Armano, Antonio Corradi, and Valeria Filippi Current Status: Paper to Publisher (6/23/98). 27 pgs. Short Code: [Armano:98a]

(C390) “Simulated Annealing Task to Processor Mapping for Domain Decomposition Methods on Distributed

Parallel Computers” Authors: C. C. Pain, C. P. E. de Oliveira, and A. J. H. Goddard Current Status: Paper to Publisher (6/23/98). 11 pgs. Short Code: [Pain:98a]

(C391) “Ordinary Linear Model Estimation on a Massively Parallel SIMD Computer”

Authors: Erricos John Kontoghiorghes Current Status: Paper to Publisher (6/23/98). 17 pgs. Short Code: [Konto:98a]